OpenStudy (jenniferjuice):

Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2. −14 −7 7 11

OpenStudy (anonymous):

Substitute x into your equation g(x) then using the value you find use it for the value of x in your f(x)

OpenStudy (jenniferjuice):

wait what do i d0?!?!

OpenStudy (kinggeorge):

When they're asking for \(f(g(x))\), they want you to first find \(g(x)\), and plug that value in for \(f(x)\). So can you tell me what \(g(2)\) is?

OpenStudy (jenniferjuice):

-7?

OpenStudy (kinggeorge):

Right. That's \(g(2)\). Now take that value, and plug it into \(f(x)\). So what's \(f(-7)\)?

OpenStudy (jenniferjuice):

-3?

OpenStudy (kinggeorge):

That's not quite right. \(f(x)\) is given by \(f(x)=-x+4\). So when you plug in \(-7\), you get\[f(-7)=-(-7)+4\]The two negatives actually cancel each other out. So you'll actually get\[f(-7)=-(-7)+4=7+4=11\]

OpenStudy (jenniferjuice):

thank you

OpenStudy (kinggeorge):

So now, if they instead asked for \(f(g(x))\) when \(x=3\) could you do that?

OpenStudy (jenniferjuice):

is it 1?

OpenStudy (kinggeorge):

That would be what you get if \(x=-3\) instead of \(3\). For positive \(3\), you would get\[g(3)=-9\]and\[f(-9)=13\]For a final result of \(13\).

OpenStudy (kinggeorge):

With \(x=-3\), you have\[g(-3)=3\]and\[f(3)=1\]

OpenStudy (jenniferjuice):

oh